For TiSO, Ksp = 3.6.10-6. What is solubility of Tl2SO4 in pure water? 1.9-10-3M 4.2-10-3M 3.6 10-4M 9.7 10-3 M
Answer #1

Answer will be 9.7 X 10-3 M.

Tl2SO4(s) \rightleftharpoons 2Tl+(aq) + SO4-2(aq)

Ksp= [Tl+][SO4-2]

If solubility of Tl2SO4 = s mol/L

Then

[Ti+]= 2s

[SO4-2] = s

Ksp= [Tl+][SO4-2] = (2s)2(s)

3.6 X10-6 = 4s3

s = 9.7 X 10-3 M